# LiDAR Remote Sensing Meets Weak Supervision: Concepts, Methods, and Perspectives

### v1

@@ Abstract
LiDAR (Light Detection and Ranging) enables rapid and accurate acquisition of three-dimensional spatial data, widely applied in remote sensing areas such as surface mapping, environmental monitoring, urban modeling, and forestry inventory. LiDAR remote sensing primarily includes data interpretation and LiDAR-based inversion. However, LiDAR interpretation typically relies on dense and precise annotations, which are costly and time-consuming. Similarly, LiDAR inversion depends on scarce supervisory signals and expensive field surveys for annotations. To address this challenge, weakly supervised learning has gained significant attention in recent years, with many methods emerging to tackle LiDAR remote sensing tasks using incomplete, inaccurate, and inexact annotations, as well as annotations from other domains. Existing review articles treat LiDAR interpretation and inversion as separate tasks. This review, for the first time, adopts a unified weakly supervised learning perspective to systematically examine research on both LiDAR interpretation and inversion. We summarize the latest advancements, provide a comprehensive review of the development and application of weakly supervised techniques in LiDAR remote sensing, and discuss potential future research directions in this field.
@@
LiDAR remote sensing , Weakly supervised learning , LiDAR Point clouds , LiDAR Waveform , Ubiquitous annotations
@@ 1 Introduction
LiDAR (Light Detection and Ranging) has emerged as a cornerstone of modern remote sensing, offering unparalleled capabilities in capturing high-precision, three-dimensional data of the Earth’s surface [1, 2]. Its unique ability to measure distances with high accuracy, penetrate tree canopy, and generate 3D point clouds has driven its adoption across a wide range of applications. From urban modeling and forestry monitoring to ecological and environmental studies, LiDAR’s versatility has transformed how we analyze and interpret the physical world. Additionally, it aids various applications, such as mapping terrain under dense vegetation, and facilitating urban management.
From an interdisciplinary perspective, LiDAR remote sensing technology plays a dual role across various application fields: it serves both as a data source that requires interpretation and as a sparse signal that supports large-scale remote sensing inversion. As a data source, LiDAR data is derived from different platforms and sensors, including space-based systems (e.g., ICESat-2), airborne systems (e.g., drones and aircraft), and ground-based systems (e.g., sensors mounted on vehicles or tripods). Its interpretation include denoising [3, 4], filtering [5], registration [6], point cloud segmentation [7, 8], and 3D reconstruction [9]. On the other hand, LiDAR, as a weak supervisory signal in remote sensing inversion, plays a crucial role, especially when combined with optical data. It is widely applied in large-scale retrieval of surface properties such as forest height, building height, biomass, leaf area index [10], sea ice thickness [11], and snow depth [12], as shown in Fig. 1.
@@
LiDAR interpretation and inversion face several common challenges. Firstly, LiDAR interpretation often relies on a large number of accurate manual annotations, but the high quality of such annotations requires significant time and resources, making the process costly in practical applications [13]. Similarly, LiDAR inversion also encounters high annotation costs. Due to the sparsity of LiDAR data, large-scale, continuous surface mapping is difficult to achieve independently [14]. However, acquiring more annotated data typically necessitates field surveys to measure tree diameter at forest height, tree height, leaf area index, etc., which is time- and resource-intensive [15]. Secondly, both LiDAR interpretation and inversion face domain shift issues [16, 17]. LiDAR data collected from different platforms exhibit significant variations in density, resolution, coverage, and noise levels, leading to notable domain shifts. Additionally, the Earth’s surface is diverse and constantly changing, covering various terrains such as cities, forests, mountains, and water bodies, which further exacerbates domain shift problems.
In response to these challenges, weakly supervised learning provides an effective solution [18]. Unlike traditional supervised learning that relies on large amounts of high-quality annotated data, weakly supervised learning combines limited annotated data with unannotated or noisy annotated data, leveraging techniques such as consistency regularization [19], self-training [20], pseudo-labeling [21], active learning [22], and feature alignment [23]. This optimizes the model’s learning process, allowing for more accurate results even in the context of scarce annotations. This approach offers a new perspective for LiDAR data interpretation and inversion and holds promise for addressing issues related to high annotation costs, domain shift, and data sparsity.
@@
Figure 1: An overview of LiDAR remote sensing meets weak supervision. LiDAR remote sensing data are primarily obtained from satellite, airborne, vehicle-mounted, and ground-based platforms, providing rich surface information. The interpretation tasks for LiDAR data include denoising, filtering, registration, semantic segmentation, instance segmentation, and 3D reconstruction. Inversion tasks involve estimating forest height, building height, biomass, leaf area index, sea ice thickness, and snow depth. These tasks face challenges such as high annotation costs, domain shift, difficulties in field validation, and sparse supervisory signals. To address these challenges, weakly supervised learning techniques such as incomplete, inaccurate, inexact supervision, and domain adaptation can be employed. Additionally, methods like consistency regularization, pseudo-labeling, self-training, active learning, and feature alignment can effectively improve model generalization. These images are obtained from [2, 4, 24, 8, 10, 12, 11, 7, 5, 6, 9, 25].
However, most researchers have focused solely on LiDAR data interpretation under weak supervision or only on LiDAR-assisted large-scale remote sensing inversion. Few studies have been able to examine LiDAR remote sensing from a unified perspective, considering the dual role of LiDAR. Although several reviews on point clouds have been published, they have certain limitations. The survey in [26] covers a wide range of topics without focusing on LiDAR remote sensing, overlooking reviews in the field of remote sensing inversion. Works like [27, 28] focus on fully supervised learning, neglecting weakly supervised approaches, and [29] provides systematic reviews only on unsupervised learning. In summary, a thorough and comprehensive review of LiDAR remote sensing from the perspective of weak supervision is highly needed and timely.
@@
To the gaps between different research directions and advancing the development of LiDAR remote sensing, this review focuses on two primary aspects: the interpretation of LiDAR remote sensing data under weak supervision and the use of LiDAR as a weak supervisory signal in remote sensing, and aims to comprehensively explore the potential and challenges of LiDAR remote sensing in various applications from the perspective of weak supervision. Fig. 2 provides an overview of the entire article. The contributions of this paper are summarized as follows:
@@
    To the best of our knowledge, this is the first review to examine LiDAR remote sensing from a unified perspective, considering its dual role as both a data source and weak supervision signals.
@@
    This paper highlights promising research directions in LiDAR remote sensing from a weak supervision perspective and explores the potential for integrating more weakly supervised learning techniques into LiDAR-aided inversion.
The structure of this paper is as follows. Section 2 reviews weak supervised interpretation of LiDAR data. Section 3 reviews that LiDAR serves as a weak supervisory signal when combined with optical data. Section 4 discusses our perspectives and reveals possible research directions in the future. Finally, Section 5 concludes the paper.
Figure 2: A taxonomy of methods and applications for LiDAR remote sensing meets weak supervision. LiDAR interpretation is primarily divided into incomplete supervision, inexact supervision, inaccurate supervision, and domain adaptation. Based on related literature, these tasks can further be categorized into semi-supervision, sparse supervision, scene-level supervision, noisy labels, unsupervised domain adaptation, and others. For LiDAR inversion, from an application perspective, they are mainly classified into canopy height, building height, and above-ground biomass estimation. Technically, inversion methods are divided into generalization approaches and the use of LiDAR as a weak supervision signal.
2 LiDAR Interpretation with Weak Supervision
@@
Weak supervision Setting Task Method Contribution
Incomplete Supervision Semi supervision Semantic segmentation SSGF[30] Handle labeled and unlabeled data via graphs separately
@@
Object detection - -
Instance segmentation - -
Sparse supervision Semantic segmentation Wang et al. [36] Select reliable pseudo-label via adaptive thresholding
@@
Object detection - -
Instance segmentation - -
@@
Subcloud level supervision Semantic segmentation Lin et al. [39] Explore semantic information between adjacent subclouds through overlap region loss
Scene level supervision Semantic segmentation Wang et al. [40] Select informative subclouds for annotation using TOD-based active learning
Scribble level supervision - - -
Inaccurate Supervision Noisy label supervision Semantic segmentation ADNLC [41] Help the network learn true labels by iteratively updating noisy labels
Domain Adaptation Unsupervised domain adaptation Semantic segmentation Luo et al. [23] Align source and target domain feature distributions using an MCD-based adversarial framework
@@
PS-UDA [46] Train the target domain GCN classifier using enhanced target domain features and high-quality pseudo-labels
Source free domain adaptation - - -
@@
Test time adaptation Semantic segmentation Yao et al. [16] Update BN stats and affine parameters using test batch statistics and a self-supervised module
In LiDAR remote sensing, we deal with massive point cloud data from various sources, such as airborne and space-borne LiDAR. High-precision interpretation of LiDAR data heavily relies on supervised machine learning models, which, in turn, require large volumes of high-quality and densely annotated data. However, acquiring such annotations is costly, particularly in the context of Earth observation, where regional variations, differences in land cover, and temporal changes in these features pose significant challenges to both the accuracy and generalization capability of models. To address this issue, weakly supervised learning methods have gained increasing attention in recent years. These approaches allow for training models with limited or imprecise labels, significantly alleviating the need for extensive annotations. As a result, weakly supervised learning has become a crucial and promising direction for LiDAR data interpretation, offering a path forward in overcoming the limitations of traditional supervision.

### v2

@@ Abstract
Light detection and ranging (LiDAR) remote sensing encompasses two major directions: data interpretation and parameter inversion. However, both directions rely heavily on costly and labor-intensive labeled data and field measurements, which constrains their scalability and spatiotemporal adaptability. Weakly Supervised Learning (WSL) provides a unified framework to address these limitations. This paper departs from the traditional view that treats interpretation and inversion as separate tasks and offers a systematic review of recent advances in LiDAR remote sensing from a unified WSL perspective. We cover typical WSL settings—including incomplete supervision (e.g., sparse point labels), inexact supervision (e.g., scene-level tags), inaccurate supervision (e.g., noisy labels), and cross-domain supervision (e.g., domain adaptation/generalization)—and corresponding techniques such as pseudo-labeling, consistency regularization, self-training, and label refinement, which collectively enable robust learning from limited and weak annotations. We further analyze LiDAR-specific challenges (e.g., irregular geometry, data sparsity, domain heterogeneity) that require tailored weak supervision, and examine how sparse LiDAR observations can guide joint learning with other remote-sensing data for continuous surface-parameter retrieval. Finally, we highlight future directions where WSL acts as a bridge between LiDAR and foundation models to leverage large-scale multimodal datasets and reduce labeling costs, while also enabling broader WSL-driven advances in generalization, open-world adaptation, and scalable LiDAR remote sensing.
@@
LiDAR remote sensing , Weakly supervised learning , Point clouds , Domain shift , Ubiquitous annotations
@@ 1 Introduction
Light detection and ranging (LiDAR), an active remote-sensing modality typically based on laser time-of-flight ranging, has emerged as a cornerstone of modern remote sensing, offering exceptional capabilities in capturing high-precision 3D data of the Earth’s surface [1, 2]. LiDAR remote sensing has reshaped geospatial analysis, supporting applications in terrain mapping, ecological monitoring, agricultural and forestry resource management, cryosphere observations, urban modeling, as well as infrastructure inspection and hazard assessment.
From an interdisciplinary perspective, LiDAR remote sensing plays a dual role across diverse applications: it serves as a data source requiring interpretation while also acting as a sparse signal supporting large-scale remote sensing inversion. As a data source, LiDAR measurements are collected from multiple platforms, typically including spaceborne systems (e.g., Global Ecosystem Dynamics Investigation (GEDI) and Ice, Cloud, and land Elevation Satellite-2 (ICESat-2)), airborne laser scanning (ALS), terrestrial laser scanning (TLS), and mobile laser scanning (MLS). These platforms collect massive volumes of high-resolution 3D data, which cannot be directly applied in downstream analyses. Instead, the raw data must undergo interpretation—typically involving preprocessing (e.g., denoising  [3, 4]), filtering  [5], classification [6], segmentation  [7], and reconstruction  [8]—to derive meaningful geometric and semantic information that supports analytical and operational applications. As a sparse signal, LiDAR provides physically grounded, high-quality observations that serve as critical constraints or weak supervision for large-scale land surface parameter inversion. Termed LiDAR-based inversion, this involves exploiting LiDAR data and machine learning to map sparse yet physically grounded observations to geophysical and geometric parameters (e.g., building height [9], canopy height [10], leaf area index [11], biomass [12], sea ice thickness [13], and snow depth [14]), thereby enabling large-area parameter retrieval from limited reference measurements, as illustrated in Figure 1.
@@
LiDAR data interpretation and LiDAR-based inversion face several common challenges. Firstly, LiDAR data interpretation often relies on a large number of accurate manual annotations, but the high quality of such annotations requires significant time and resources, making the process costly in practical applications [15]. Similarly, LiDAR-based inversion also encounters high annotation costs. Due to the sparsity of LiDAR data, large-scale, continuous surface mapping is difficult to achieve independently [16]. However, acquiring more annotated data typically necessitates field surveys to measure diameter at breast height, tree height, leaf area index, etc., which is time- and resource-intensive [17]. Secondly, both LiDAR data interpretation and LiDAR-based inversion face domain shift issues [18, 19]. LiDAR data collected from different platforms exhibit significant variations in density, resolution, coverage, and noise levels, leading to notable domain shifts. Additionally, the Earth’s surface is diverse and constantly changing, covering various terrains such as cities, forests, mountains, and water bodies, which further exacerbates domain shift problems.
In response to these challenges, weakly supervised learning (WSL) provides an effective solution [20]. WSL is a broad machine learning paradigm that constructs predictive models under limited, weak, or mismatched supervision, encompassing settings with incomplete, inexact, or noisy labels as well as cross-domain scenarios such as domain adaptation and domain generalization. Unlike fully supervised learning, which relies on large amounts of precise annotations, weakly supervised approaches can leverage sparse, uncertain, or heterogeneous supervision (e.g., crowd-sourced annotations) to substantially reduce annotation costs while maintaining generalization. Building on this extended taxonomy, We adapt WSL as a machine learning paradigm for LiDAR remote sensing, in which models leverage sparse, coarse, noisy, or domain-shifted supervisory signals—together with intrinsic data structures and explicit priors—to learn representations and inference mechanisms that support dense and accurate data interpretation and inversion. In the context of LiDAR data interpretation, weak supervision enables the extraction of meaningful geometric and semantic information (e.g., structural modeling and semantic segmentation) from limited labels through techniques such as consistency regularization [21], self-training [22], and pseudo-labeling [23]. For LiDAR-based inversion, we treat LiDAR measurements as sparse yet physically grounded supervision that anchors large-scale parameter retrieval: models fuse dense imagery with these incomplete LiDAR observations to learn a sparse-to-dense mapping from observations to geophysical and biophysical parameters.
@@
Figure 1: An overview of LiDAR remote sensing meets weak supervision. LiDAR remote sensing data are primarily obtained from satellite, airborne, vehicle-mounted, and ground-based platforms, providing rich surface information. The interpretation tasks for LiDAR data include denoising, filtering, registration, semantic segmentation, instance segmentation, and 3D reconstruction. Inversion tasks involve estimating forest height, building height, biomass, leaf area index, sea ice thickness, and snow depth. These tasks face challenges such as high annotation costs, domain shift, difficulties in field validation, and sparse supervisory signals. To address these challenges, weakly supervised learning techniques such as incomplete, inaccurate, inexact supervision, domain generalization, and domain adaptation can be employed. Additionally, methods like consistency regularization, pseudo-labeling, self-training, active learning, and feature alignment can effectively improve model generalization. These images are obtained from [2, 4, 24, 25, 11, 14, 13, 26, 5, 27, 8, 28].
However, most researchers have focused solely on LiDAR data interpretation under weak supervision or only on LiDAR-assisted large-scale remote sensing inversion. Few studies have been able to examine LiDAR remote sensing from a unified perspective, considering the dual role of LiDAR. Although several reviews on point clouds have been published, they have certain limitations. The survey [29] systematically examines key methodologies such as data augmentation, domain adaptation, WSL, and pre-trained foundation models, discussing research progress in tasks like object detection, classification, and segmentation across synthetic datasets, indoor scenes, and autonomous driving scenes. Notably, the scope of this review excludes remote sensing data (e.g., ALS and spaceborne photon-counting point clouds) and remote sensing inversion tasks, such as the estimation of canopy and building heights. Guo et al. [30] and He et al. [31] provide a comprehensive review for 3D point-cloud processing, covering classification, detection, and segmentation. However, their discussion of techniques is largely confined to fully supervised methods and overlooks the potential of WSL. Moreover, their review omits the increasingly important area of remote-sensing LiDAR point clouds. Xiao et al. [32] systematically reviews various unsupervised deep learning methods that learn general and robust feature representations by performing pretext tasks, such as generative or contrastive learning, on large-scale unlabeled 3D point clouds. These learned representations can then be effectively transferred to downstream supervised tasks. However, the survey does not sufficiently discuss the role and potential of WSL in LiDAR remote sensing. Therefore, a systematic and comprehensive review of LiDAR remote sensing from a weakly supervised perspective is both necessary and timely.
@@
To bridge the gaps between different research directions and advance the development of LiDAR remote sensing, this review focuses on two primary aspects: the interpretation of LiDAR remote sensing data under weak supervision and the use of LiDAR as a weak supervisory signal in remote sensing. It aims to comprehensively explore the potential and challenges of LiDAR remote sensing in various applications from the perspective of weak supervision. Fig. 2 provides an overview of the entire article. The contributions of this paper are summarized as follows:
@@
    To the best of our knowledge, this is the first review to examine LiDAR remote sensing from a unified perspective, considering its dual role as both a data source and a weak supervisory signal.
@@
    This paper highlights promising research directions in LiDAR remote sensing from a weak supervision perspective and explores the potential for integrating more weakly supervised learning techniques into LiDAR-based inversion.
This paper is organized as follows. Section 2 provides the background, outlining specialized deep learning architectures for LiDAR point clouds and the rationale for the paradigm shift from fully to weakly supervised learning. Sections 3 and 4 form the core of the paper, reviewing the applications of weakly supervised learning in LiDAR data interpretation and parameter inversion, respectively. Section 5 offers a forward-looking perspective, analyzing the unique characteristics of LiDAR data, the limitations of current methods, the potential of foundation models, and future research directions. Finally, Section 6 concludes the paper.
Figure 2: A taxonomy of methods and applications for LiDAR remote sensing meets weak supervision. Weakly supervised approaches for LiDAR remote sensing include incomplete supervision, inexact supervision, inaccurate supervision, domain adaptation and domain generalization. These approaches jointly support point cloud interpretation and LiDAR-based inversion tasks.
2 Preliminaries
2.1 Deep learning for LiDAR
LiDAR data primarily include point clouds acquired from ground-based, airborne, and spaceborne platforms, as well as large-footprint full-waveform measurements. To process the massive amount of LiDAR data, deep learning–based methods are commonly employed to extract semantic and geometric information for downstream applications. However, point cloud data are inherently unstructured, unordered, and distributed in three-dimensional space, which makes conventional image-based neural network architectures inapplicable. This stems from the fact that these architectures—such as convolutional and transformer-based networks—rely on regular grid structures to enable local feature aggregation and position encoding. In contrast, point clouds lack such fixed adjacency and ordering, making it difficult to directly apply conventional deep learning architectures designed for regular grids.
Therefore, to accommodate the unstructured, unordered, and highly sparse nature of point clouds, researchers have developed network families tailored to 3D geometry rather than regular image grids. Point-based models (e.g., PointNet [33]/PointNet++ [34]) operate directly on 3D coordinates with permutation-invariant set functions and local neighborhood aggregation; graph-based models (e.g., DGCNN [35]) build dynamic kNN/radius graphs to capture adaptive local topology; kernel/stateful convolutions and efficient sampling (e.g., KPConv [36], RandLA-Net [37]) improve locality modeling and scalability on large scenes. In parallel, voxel-based pipelines (e.g., VoxelNet [38], submanifold/Minkowski sparse 3D CNNs [39]) discretize space and leverage sparse tensors for efficient convolutions. More recently, Transformer-style architectures (e.g., Point Transformer [40] and its variants) employ self-attention to model long-range dependencies and cross-scale interactions, often combined with local grouping or sparse attention for efficiency. Most of these architectures operate directly on irregular 3D points, explicitly handling nonuniform density, occlusions, and sparsity to achieve efficient and robust 3D representation learning. For a detailed review for 3D deep learning, we refer the reader to [41] and [42].
2.2 From fully to weakly supervised learning
With the rapid development of neural architectures for LiDAR data processing, training robust, accurate, and generalizable deep learning models for point cloud data has become a central challenge. Most existing approaches rely on fully supervised learning, a paradigm in which models are trained on densely labeled datasets—each 3D point is assigned a ground-truth class or attribute, and the learning objective minimizes the discrepancy between predictions and reference labels. This setting enables direct end-to-end optimization and often achieves high predictive accuracy. Fully supervised learning has led to significant advances in 3D scene understanding. For example, RandLA-Net achieved 83.1% overall accuracy (OA) and 70.1% mF1 on ISPRS; KPConv reported over 81% mIoU on DALES.
However, unlike 2D image tasks—where annotations can often be obtained at low cost from web data or automatically generated using vision-language models—fully supervised training for LiDAR data faces prohibitively high labeling costs. For point cloud interpretation, annotations rely heavily on manual labeling with specialized 3D software, which is labor-intensive, error-prone, and difficult to scale. For LiDAR-based inversion, collecting ground reference data is expensive, requires intensive fieldwork, and may pose safety risks. Moreover, LiDAR remote sensing targets the Earth’s surface, which is vast, heterogeneous, and continuously changing due to seasonal dynamics, human activities, and natural disturbances. These factors make it extremely difficult to obtain high-quality, spatiotemporally comprehensive annotations, rendering fully supervised training unsustainable at scale.
To tackle these challenges, weakly supervised learning (WSL) has emerged as a promising paradigm in LiDAR remote sensing. Instead of relying on exhaustive point-level annotations, WSL leverages limited or imprecise labels to train deep learning architectures specialized for point clouds, thereby substantially reducing the dependence on large-scale manual labeling in data interpretation. In the context of LiDAR-based inversion, WSL enables the use of sparse field measurements collected over limited areas to support large-scale parameter retrieval. By incorporating incomplete supervision and promoting domain generalization, weak supervision can mitigate the effects of spatial heterogeneity and temporal variability, allowing models trained on sparse local data to generalize effectively across broader regions. Consequently, WSL provides a practical and scalable alternative to conventional supervised paradigms and represents a critical research direction for advancing both LiDAR data interpretation and parameter inversion.
Table 1: A summary of commonly used semantic segmentation datasets for LiDAR remote sensing
Dataset Platform Year Area/Length # Classes # Points
Semantic3D [43] Terrestrial Laser Scanning 2017 - 8 4000M
Okland [44] Mobile Laser Scanning 2009 $1.5\times 10^{3}\>m$ 5 1.6M
Paris-rue-Madame [45] 2014 $0.16\times 10^{3}\>m$ 26 20M
iQmulus [46] 2015 $0.2\times 10^{3}\>m$ 22 12M
TUM-City-campus [47] 2016 $0.97\times 10^{3}\>m$ 8 41M
Paris-Lille-3D [48] 2018 $1.94\times 10^{3}\>m$ 9 143M
Toronto-3D [49] 2020 $1\times 10^{3}\>m$ 8 78.3M
WHU-Urban3D [25] 2023 $6.5\times 10^{3}\>m$ 18(35) 393M
ISPRS [50] Aerial Laser Scanning 2012 - 9 1.2M
DublinCity [51] 2019 $2\times 10^{6}\>m^{2}$ 13 260M
DALES [26] 2020 $10\times 10^{6}\>m^{2}$ 8 505M
LASDU [52] 2020 $1.02\times 10^{6}\>m^{2}$ 5 3.12M
Hessigheim 3D [24] 2021 $0.19\times 10^{6}\>m^{2}$ 11 125.7M/36.76M
WHU-Urban3D [25] 2023 $3.2\times 10^{6}\>m^{2}$ 9 213M
ECLAIR [53] 2024 $0.3\times 10^{6}\>m^{2} 11$ 582M
FRACTAL [54] 2024 $250\times 10^{6}\>m^{2}$ 7 9261M
@@ 2.2 From fully to weakly supervised learning
Weak supervision Setting Method Contribution
Incomplete Supervision Semi supervision SSGF[55] Handle labeled and unlabeled data via graphs separately
@@
Sparse supervision Wang et al. [60] Select reliable pseudo-label via adaptive thresholding
@@
PSD [61] Establish point cloud graph topology to enable information propagation between labeled and unlabeled points
Zhang et al. [62] Transfer prior knowledge from unlabeled data to weakly supervised networks via self-supervised point cloud colorization pre-training
SQN [63] Enhance sparse (0.1%) annotations using semantic homogeneity in point cloud neighborhoods
@@
MSCs[65] Apply scene-level constraints at the encoder, decoder, and classifier stages to alleviate class imbalance
Chen et al. [66] Use line and plane points as additional supervisory signals for hierarchical feature learning
FPWS-Net [67] Enforce consistency on unlabeled data via mutual pseudo-labeling loss
SCSQ-Net [68] Share weights between encoder and semantic inference to efficiently integrate neighborhood context
Li et al. [69] Decouples feature learning and the classifier via alternating optimization
HSCN [70] Enforce global distribution alignment with GAL and multiscale neighborhood-structure alignment with similarity metric loss
DR-Net [71] Enhance feature representation by aggregating local features of the point cloud
DSDCL [72] Employ context-awareness to enhance the model’s ability to differentiate similar points
DCUF-Net [73] Use an uncertainty-aware loss to extract reliable supervisory signals from large unlabeled datasets
Active learning Polewski et al. [74] Combine semi-supervised and active learning under Rényi entropy regularization
Luo et al. [75] Employ an active learning strategy to iteratively annotate a limited number of supervoxels
Lin et al. [76] Propose a segmentation-based active learning strategy that iteratively annotates informative unlabeled samples
Lin et al. [77] Use point entropy, segmentation entropy, and mutual information to iteratively select the most informative samples
Kölle et al. [78] Combine active learning with crowdsourced annotation
SSDR-AL [79] Select the most informative and representative samples through a graph reasoning network
Kölle et al. [80] Validate Active Learning performance
Wang et al. [81] Integrate the novel ’One-Class-One-Click’ method with uncertainty-driven active learning
Dai et al. [82] Introduce an Active Learning strategy with a Spatially Incremental Sampling method for label acquisition
@@
Few-shot learning Thr-MPRNet [84] Construct a 3D few-shot framework with feature and relation learners
Huang et al. [85] Combine few-shot learning with distribution calibration
Zhang et al. [86] Introduce Pole-NN, a training-free, non-parametric feature extractor
Dai et al. [87] Decompose cross-domain ALS segmentation into transfer learning and class-incremental learning
Yuan et al. [88] Propose a Multi-branch Feature Extractor with three-stage transfer learning: pre-training, cross-domain alignment, and incremental learning
@@
Subcloud level supervision Lin et al. [89] Explore semantic information between adjacent subclouds through overlap region loss
Scene level supervision Wang et al. [81] Select informative subclouds for annotation using TOD-based active learning
WHCN [90] Construct a hypergraph from superpoints to generate point-level pseudo-labels
Scribble level supervision - -
Inaccurate Supervision Noisy label supervision ADNLC [91] Help the network learn true labels by iteratively updating noisy labels
Domain Adaptation Unsupervised domain adaptation Luo et al. [92] Align source and target domain feature distributions using an MCD-based adversarial framework
@@
Wang et al. [97] Use geometric transformation augmentation to enable cross-modal learning from LiDAR to SAR point clouds
PS-UDA [98] Employ an L2-paradigm constraint and a Laplace matrix to pre-align cross-scene features
Luo et al. [99] Reconstruct MLS scene surfaces and align source and target point clouds in a canonical domain at the point level
Source-free domain adaptation - -
@@
Test-time adaptation Yao et al. [18] Update BN stats and affine parameters using test batch statistics and a self-supervised module
APCoTTA [100] Update partial layers and interpolating parameters to retain source knowledge
Table 3: Weakly supervised LiDAR remote sensing interpretation results on different datasets.(“1” denotes only one point; “1%” indicates that 1% of points are randomly labeled for each class. “AL” stands for active learning, and “FW” denotes few-shot learning. )
Method Backbone Supervison ISPRS LASDU H3D DALES Toronto3D WHU-MLS Semantic3D
OA mIoU mF1 OA mIoU mF1 OA mIoU mF1 OA mIoU mF1 OA mIoU mF1 OA mIoU mF1 OA mIoU mF1
PointNet++ [34] 100% 81.2 65.6 95.7 68.3 84.88 41.81
PointNet++(MSG) [34] 100% 92.56 59.47
DGCNN [35] 100% 94.24 61.79
RandLA-Net [37] 100% 83.1 70.1 97.1 80.0 95.63 77.72 94.8 77.4
KPConv [36] 100% 84.2 70.5 97.8 81.1 95.39 69.11 92.9 74.6
ShellNet [101] 100% 93.2 69.3
PointGCR [102] 100% 92.1 69.5
Wang et al. [60] KPConv 15 80.4 63.1
30 81.1 65.4
60 83.0 68.6
100 83.7 70.2
Wang et al. [20] KPConv 0.01% 85.8 72.1
0.05% 78.2 61.9 87.3 74.6 86.7 75.0
0.1% 83.0 70.0 87.6 77.4 88.2 76.3
0.2% 83.4 70.7 87.7 78.1 87.5 76.6
0.5% 84.0 71.9 87.6 78.3
PSD [61] RandLA-Net 1% 96.2 76.2 94.3 75.8
Zhang et al. [62] RandLA-Net 1% 93.7 72.6
10% 94.0 73.3
SQN [63] RandLA-Net 0.01% 95.9 60.4 94.2 68.2 90.3 65.6
0.1% 97.0 72.0 96.7 77.7 93.7 74.7
1% 96.9 80.2
WSPointNet [64] RandLA-Net 0.1% 96.76 78.96 90.83 56.48
MSCs[65] RandLA-Net 1% 96.8 81.2
Chen et al. [66] KPConv 0.05% 73.5 55.1 83.2 71.6
0.1% 78.2 65.6 84.6 74.5
0.5% 82.3 70.1 84.9 75.3
Li et al. [69] RandLA-Net 1pt 92.8 66.2
1% 94.9 76.9
HSCN [70] KPConv 0.1% 79.5 68.7 85.2 75.8
1% 82.9 72.8 86.5 77.7
DR-Net [71] SQN [63] 0.1% 96.6 72.9 96.9 78.1
1% 97.0 78.0 96.6 79.1
DSDCL [72] RandLA-Net 0.1% 79.7 65.7 85.4 73.4
1% 81.8 70.1 87.3 77.0
DCUF-Net [73] RandLA-Net 1% 97.2 81.6
Kölle et al. [80] Random Forest AL 81.59 67.46 84.84 70.39
Submanifold Sparse CNN AL 80.91 66.20 85.44 70.90
Wang et al. [81] KPConv 750(subcloud) 86.46 74.19
1320(subcloud) 94.61 85.20
Dai et al. [82] SPVCNN [103] AL 70.41
DAAL-WS [83] RandLA-Net 0.015% 97.60 81.91
0.03% 60.36
Huang et al. [85] PointSIFT [104] TUM-MLS-2016(FW) 81.2 37.5 45.1
DALES(FW) 80.7 55.9 65.0
Thr-MPRNet [84] DGCNN SensatUrban(FW) 41.40 31.58 44.06 32.41 76.82 65.25
Dai et al. [87] DGCNN SensatUrban(FW) 45.34 31.85 48.11 37.27 82.51 76.78
Yuan et al. [88] Point Transformer [40] DALES and Dublin(FW) 0.1% 81.8 65.1
1% 83.7 68.4
10% 85.5 74.0
100% 85.7 74.7
SensatUrban(FW) 84.2 58.1
Lin et al. [89] KPConv 0.5%(subcloud) 80.0 63.1
1%(subcloud) 81.4 65.9
WHCN [90] KPConv scene 86.2 58.3
ADNLC [91] KPConv noisy lables 78.2 53.1 80.5 53.5
3 LiDAR Interpretation with Weak Supervision